### v1

# Test-Time Scaling with Reflective Generative Model

Abstract
We introduce our first reflective generative model MetaStone-S1, which obtains OpenAI o3’s performance via the self-supervised process reward model (SPRM). Through sharing the backbone network and using task-specific heads for next token prediction and process scoring respectively, SPRM successfully integrates the policy model and process reward model(PRM) into a unified interface without extra process annotation, reducing over 99% PRM parameters for efficient reasoning. Equipped with SPRM, MetaStone-S1 is naturally suitable for test time scaling (TTS), and we provide three reasoning effort modes (low, medium, and high), based on the controllable thinking length. Moreover, we empirically establish a scaling law that reveals the relationship between total thinking computation and TTS performance. Experiments demonstrate that our MetaStone-S1 achieves comparable performance to OpenAI-o3-mini’s series with only 32B parameter size. To support the research community, we have open-sourced MetaStone-S1 at https://github.com/MetaStone-AI/MetaStone-S1.
Figure 1: Benchmark performance of MetaStone-S1.
1 Introduction
Over the past two years, the field of Large Language Models (LLMs) has experienced rapid advancements, marked by the emergence of increasingly sophisticated models. Notable developments include OpenAI’s GPT-4, Google’s Gemini, Meta’s LLaMA series, Alibaba’s Qwen, and DeepSeek’s R1, which have collectively pushed the boundaries of natural language understanding and generation. This progress is attributed to innovations in model architectures and training techniques, enabling LLMs to process and generate content across various formats.
Recent analyses suggest that OpenAI’s o3 model achieves its advanced reasoning and coding capabilities through Test-Time Scaling (TTS) techniques such as massive sampling, candidate scoring, and search over multiple reasoning paths (Labs, 2025; Zeff, 2024). For instance, during ARC-AGI and competitive coding evaluations, o3 was shown to generate up to 1024 candidate samples for each query (Chollet, 2024; OpenAI, 2025). These inference-time strategies mark a significant shift from traditional one-pass models, enabling o3 to adapt dynamically to novel tasks and achieve near-human performance in reasoning benchmarks.
TTS approaches can be categorized into two types: internal TTS and external TTS. Internal TTS strategies use CoT for longer thinking processes (Guo et al., 2025; OpenAI, 2024), which benefits from Long-CoT Supervised Fine-Tuning and reinforcement learning. Recent internal TTS methods (Guo et al., 2025) mainly suffer from the false positive reasoning process, as the outcome reward will misclassify the correct answer with incorrect reasoning during the training stage. External TTS is proposed for selecting correct reasoning process, which is proved to be more effective in performance boosting compared with outcome reward (Lightman et al., 2023). Prominent external TTS algorithms include Best-of-N sampling, Beam Search, and Diverse Verifier Tree Search, using the Process Reward Model(PRM) to select the sampling and search-based candidates. For example, Best-of-N firstly generates multiple outputs, and then sequentially uses PRM-based to score and select the best solution. Liu et al. (2025) point out that the best external TTS solution varies from policy model with different parameters. When the parameter of the policy model is less than 32B, the search methods (Beam Search and Diverse Verifier Tree Search) achieve better results than the sampling one(Best-of-N). However, when the parameter size is equal to or greater than 32B, the sampling method can achieve better performance. Based on the above analysis, internal and external TTS are two individual methods and can benefit each other, e.g. using the Best-of-N to boost the Long-CoT model with PRM.
This paper focuses on external TTS and proposes a Reflective Generative Form that enables LLMs to integrate reasoning and evaluation within a single model. We first design a Self-supervised Process Reward Model (SPRM), which shares its backbone network with the policy model. By employing task-specific heads for next-token prediction and process scoring, SPRM effectively unifies the policy model and process reward model(PRM) into a unified interface. Compared to previous PRMs, this unified design provides a more efficient scoring process with little parameter and computational overhead. To reduce the need for expensive process annotations, we further introduce a Self-supervised Process Reward Loss, which optimizes the SPRM using only final answers. Based on SPRM, we propose MetaStone-S1, containing high, medium, and low reasoning modes, with the controllable thinking length. Experiment results show that MetaStone-S1 achieves comparable performance to o3-mini’s series with only 32B parameters.
1.1 Contributions
Reflective generative form: A unified interface for policy model and process reward model
  • We share the backbone parameters between the policy model and SPRM, and design two task-specific heads for token prediction and process scoring respectively (shown in Fig.2). As the policy model has been trained on T-level data during the pre-training stage, the unified interface provides a stable cold-start weight for SPRM, and provides a high way to share knowledge between the two models.
  • The unified interface enables the single network with both reasoning and scoring capabilities. Benefiting from the concise design of the SPRM head, we significantly reduce the extra parameters of the reward model by at least 99%.
  • The unified interface is naturally suitable for TTS. We show a scaling law that reveals the relationship between total thinking computation and TTS performance. Finally, MetaStone-S1 achieves comparable performance as the o3-mini’s series with only 32B parameter size.
Self-superviesd process reward model: Aligning the label of PRM to the outcome reward
[…]
  • The SPRM has a large potential for scaling-up: 1)We find the aha-moment of the SPRM, where the model significantly improves the discrimination between positive and negative thinking processes at a certain checkpoint. 2) SPRM has strong robustness, which is trained on only mathematical data and provides significant improvement in mathematical, coding, and scientific performance.
2 Related Works
2.1 Test-Time Scaling
Test-Time Scaling (TTS) is a technique that leverages additional computational resources at inference time to tackle challenging problems. With the remarkable performance improvements demonstrated by OpenAI o1 (Jaech et al., 2024), TTS has become a research hot spot for enhancing the reasoning capabilities of LLMs. TTS can be divided into two categories: internal TTS and external TTS. Internal TTS introduces the long Chain-of-Thought (CoT) to generate answers based on the detailed reasoning process. OpenAI o1(Jaech et al., 2024) and DeepSeek R1(Guo et al., 2025) introduce a thinking process to plan the solution and guide the final answer. Jin et al. (2024); Yeo et al. (2025) have shown that long CoT can help models correct mistakes by themselves and decompose complex problems more effectively, thereby improving performance. DeepScaler(Luo et al., 2025) demonstrates that by carefully extending the context length during training, only a 1.5B-parameter model can surpass the O1-Preview. However, Chen et al. (2024b, a) have highlighted the risk of overthinking, where excessively long reasoning trajectories may lead to performance degradation. On the other hand, external TTS scaling up inference through search-based strategies and auxiliary reward models. A common approach is the Best-of-N strategy (Lightman et al., 2023; Brown et al., 2024; Wang et al., 2023), which generates multiple candidates and selects the best one based on scores from the pretrained reward model. Moreover, fine-grained methods have also been explored, including, such as Beam Search (Liu et al., 2025; Snell et al., 2024), Diverse Verifier Tree Search (Beeching et al., ) and Monte Carlo Tree Search (MCTS) (Zhang et al., 2024; Guan et al., 2025; Luo et al., 2024). These methods search at the step level and utilize Process Reward Models (PRMs) to guide the reasoning trajectory step-by-step. Beyond search strategies, recent work emphasizes that the quality of the reward model is a crucial factor in external TTS (Guan et al., 2025). A straightforward and effective way to enhance a model’s reasoning ability is to develop a high-quality reward model.
2.2 Process Reward Model
Process Reward Models (PRMs) focus on evaluating LLMs at the step level. Lightman et al. (2023) unveil that this fine-grained guidance can lead to better TTS performance compared with the global-level Outcome Reward Model (ORM). However, accurately identifying logical errors in LLM outputs remains challenging, PRMs require high-quality task-specific annotated data for training. To this end, recent works Wang et al. (2023) leverage Monte Carlo estimation to automatically assign step-level scores using only the final answers as supervision. Zhang et al. (2024); Guan et al. (2025) iteratively synthesizes data by MCTS and fine-tuning both LLMs and PRMs, improving performance across both models. Tan et al. (2025) follow the LLM-as-a-judge method and introduce a new LLM to annotate the reward of each step. Nonetheless, Zhang et al. (2025) point out that labels generated by Monte Carlo estimation can be noisy, as incorrect reasoning processes may still yield correct final answers. They further propose a hybrid approach that combines both Monte Carlo estimation with the LLM-as-a-judge.
Despite these advances, existing PRMs still suffer from several challenges. First, the PRMs are trained on a new large-scale LLM model, resulting in significant training and inference costs. Second, most PRM training methods typically follow an off-policy strategy, which limits their ability to directly discriminate outputs generated by the target LLM. The unseen distribution during inference time may further degrade performance. To address these issues, we propose a Reflective Generative Form, which shares most parameters with the target LLM and supports on-policy optimization, enabling more efficient and aligned training.
Figure 2: The training and inference framework of Reflective Generative Models.
3 Approach
3.1 Overview
Recent Test-Time Scaling (TTS) approaches heavily rely on additional large-scale pretrained reward models (e.g., 7B-72B LLM-based PRMs Liu et al. (2025)), leading to a complex training pipeline and huge extra inference costs. In this work, we propose a Reflective Generative Form, which unifies the reasoning and evaluation processes in a single model. Specifically, a Self-supervised Process Reward Model (SPRM) is introduced, which shares most of its parameters and forward computation with the policy model. During training, we design a Self-supervised Process Reward loss to optimize SPRM only using final answer. The unified model structure and label format allow both reasoning and evaluation tasks to be trained simultaneously in a fully end-to-end manner, eliminating the need for expensive process reward labeling and complex PRM-specific training. At inference time, the policy model and SPRM share the feature extraction parameters, enabling parallel generation of solutions and corresponding evaluation scores. Compared to existing methods with separate PRMs, this design significantly improves TTS efficiency.
In the following sections, we first describe the Reflective Generative Form with SPRM. Then, we introduce the optimization process and present the Self-supervised Process Reward Loss (SPR Loss) for SPRM. Finally, we show the TTS process of Reflective Generative models.
3.2 Reflective Generative Form
The Self-supervised Process Reward Model (SPRM) is proposed to train the policy model to distinguish the solution correctness by itself. The policy model is based on the reasoning LLMs with the thinking process (denoted by the ’<think>’ and ’</think>’ tokens). The joint training framework of the policy model with SPRM is shown in Fig.2(a). In contrast to other PRM methods, SPRM does not need to add additional special tokens for segmenting steps or fine-tune the LLM for step-level output format. It can be simply cold-started from any open-source reasoning model.
Step segmentation.
[…]
Process Score Prediction.
A lightweight binary classifier (named SPRM head) is used to predict the score for each step. The SPRM head consists of two linear layers and a dropout layer. Since step-tokens represent the end of each step, we extract features from the second-to-last layer of the model at all step-token positions and feed them into the SPRM head to predict the corresponding process scores. During inference, given a thinking process with multiple steps, we calculate the geometric mean score of all steps as the final score for it, which is formulated as follows:
$\text{S}_{\text{final}}=\left(\prod_{i=1}^{n}\text{Score}_{i}\right)^{\frac{1}{n}}=\left(\prod_{i=1}^{n}SPRM(f_{token_{i}})\right)^{\frac{1}{n}},$ (1)
where $n$ is the number of steps in the given process, and $f_{token_{i}}$ denotes the feature of the $i$-th step-token from the policy model.
With SPRM, the policy model and the reward model are unified into a single interface, allowing them to be trained simultaneously. This combination presents a straightforward and simple training pipeline for on-policy PRM learning, where both models continuously refine their parameters through shared experience, ensuring consistent improvement in the quality of the generated solutions.
3.3 Self-supervised Optimization
For optimizing the policy model, we adopt Group Relative Policy Optimization (GRPO) following Shao et al. (2024). To optimize the SPRM head, we propose a Self-supervised Process Reward Loss (SPR Loss), which enables learning process discrimination ability only from final answer labels. The SPR Loss is formulated as follows,
$\mathcal{L}_{\text{SPR}}=\frac{1}{N}\sum_{i=1}^{N}w_{i}*BCELoss(\text{Score}_{i},y_{i}),\quad\text{where}\ w_{i}=\begin{cases}1,&\text{if}\ y_{i}=1\ \&\ \text{Score}_{i}>0.5\\ 1,&\text{if}\ y_{i}=0\ \&\ \text{Score}_{i}<0.5\\ 0,&\text{others}\end{cases},$ (2)
where $i$ denotes the step-tokens, $w_{i}$ is a token-level weight, $\text{Score}_{i}$ is the corresponding process score predictions, and $y_{i}$ denotes whether the final prediction is correct. In Eq.2, the process score is optimized based on the correctness of the final answer. However, since a correct final answer may include incorrect intermediate steps and vice versa (Lightman et al., 2023), we introduce the self-supervised dynamic weight $w_{i}$ to mitigate supervision noise. Specifically, we use the SPRM head’s own prediction on each step as the pseudo label and set $w_{i}=1$ only if the pseudo label is consistent with the final answer’s correctness. This dynamic filtering allows the model to avoid noisy samples and focus on the most representative steps of correct and incorrect solutions. Thus, by enlarging the score gap between correct and incorrect steps, SPRM can progressively learn the process evaluation ability with only final annotation.
3.4 Inference with Test-Time Scaling
In the inference stage, the SPRM can guide the thinking process of the policy model through TTS. The total inference stage of SPRM divides into three steps(shown in Fig.2(b)): (1) For the given question, the policy model first samples $k$ candidate thinking process: $think_{1},think_{2},\dots,think_{k}$. (2) The SPRM evaluates the steps in each process and obtains the final score by the geometric mean score of all steps in the corresponding process: $S_{1},S_{2},\dots,S_{k}$. (3) The process with the highest final score is chosen and guides the policy model to answer the question (Eq.3).
$\text{Answer}=\text{LLM}(think_{i^{*}}),\ \text{where}\ i^{*}=argmax(S_{1},S_{2},\dots,S_{k})$ (3)
4 Experiment
4.1 Baseline&Dataset
We conduct experiments on the models with three different sizes: MetaStone-S1-1.5B, 7B, and 32B, which are initialized from DeepSeek-R1-Distill-Qwen-1.5B/7B (Guo et al., 2025), and QWQ-32B (Team, 2025) with continual reinforcement training. Our training dataset is constructed from multiple publicly available math-related sources, including NuminaMath (Li et al., 2024), OpenR1-Math-220k, DeepScaleR (Luo et al., 2025), LIMR (Li et al., 2025), and OREAL-RL (Lyu et al., 2025). We apply a multi-agent data cleaning framework to ensure data quality, resulting in a final dataset of 40k high-quality examples. In the inference stage, we set 3 reasoning efforts with $k=2,8,32$ in Eq.3, named MetaStone-S1-low, -medium, and -high.
We evaluate our models on mathematical benchmarks: AIME2024 and AIME2025 (AIME, 2025) , and out-of-distribution benchmarks: LivecodeBench(240801-250201) (Jain et al., 2024) and C-Eval (Huang et al., 2023). AIME2024 and AIME2025 are challenging mathematical benchmarks, designed to assess the mathematical reasoning capabilities of LLMs. LivecodeBench(240801-250201) contains high-quality programming problems from coding competition websites (LeetCode, AtCoder, and CodeForces), designed to assess the code generation and problem-solving abilities of LLMs. C-EVAL is a comprehensive Chinese evaluation benchmark designed to assess the Chinese knowledge and reasoning abilities of LLMs. We adopt Pass@1 as the evaluation metric. For each problem, the model generates only one final answer, and the Pass@1 score is computed as the proportion of correctly solved problems. To improve the stability of the results, we repeat the evaluation 64 times and report the average accuracy as the final score.
Model Mathematical Out-of-Distribution
AIME24 AIME25 LiveCodeBench C-EVAL
Small-size Open-Sourced Models
DeepScaleR-1.5B-Preview 43.1 30.0 - -
R1-Distill-Qwen-1.5B 28.9 22.8 16.9 27.1
R1-Distill-Qwen-7B 55.5 - 37.6 -
R1-Distill-Llama-8B 50.4 - 39.6 -
Baseline-1.5B 39.3 29.9 22.4 41.8
MetaStone-S1-1.5B-low 44.0 32.6 24.2 43.6
MetaStone-S1-1.5B-medium 53.1 35.7 26.6 43.9
MetaStone-S1-1.5B-high 57.9 40.4 28.1 44.1
Baseline-7B 54.7 41.2 39.4 51.3
MetaStone-S1-7B-low 60.7 45.4 41.7 55.1
MetaStone-S1-7B-medium 66.3 48.3 44.1 57.5
MetaStone-S1-7B-high 70.2 48.6 44.4 57.8
Large-size Open-Sourced Models
s1-32B 56.7 50.0 - -
QwQ-32B 79.5 69.5 63.4 88.4
R1-Distill-Qwen-32B 72.6 49.6 57.2 82.2
GLM-Z1-32B-0414 80.8 63.6 59.1 -
DeepSeek-R1-671B 79.8 70.0 65.9 91.8
Closed-Sourced Models
Claude-3.5-Sonnet1022 16.0 7.4 37.2 76.7
GPT-4o-0513 9.3 11.6 32.9 -
OpenAI-o3-mini* 79.6 74.8 67.4 75.9
Baseline-32B 79.9 70.5 63.4 89.4
MetaStone-S1-32B-low 82.0 72.0 63.8 89.5
MetaStone-S1-32B-medium 84.2 73.4 64.0 89.6
MetaStone-S1-32B-high 85.2 73.6 64.2 89.7
Table 1: Comparison of MetaStone-S1 models and other comparable models. * denotes medium level of OpenAI-o3-mini. The best and second-best results are shown in bold and underlined.
4.2 Main Results
Table 1 summarizes the performance of our MetaStone-S1 models across four representative benchmarks. We denote baseline as the models without SPRM. Across different model scales, our proposed SPRM strategy consistently enhances the baseline, particularly on mathematical reasoning benchmarks. Specifically, compared with the baseline, MetaStone-S1 achieves performance gains of 18.6/15.5/5.3 points on AIME24 and 10.5/7.4/3.1 points on AIME25 for the 1.5B/7B/32B sizes, respectively. For other tasks, SPRM continues to offer stable improvements, yielding gains of 5.7/5.0/0.8 points on LiveCodeBench and 2.3/6.5/0.3 points on C-EVAL.
We further compare MetaStone-S1 against both advanced open-source and closed-source models. Among the open-source models, we consider DeepScaleR-1.5B-Preview (Luo et al., 2025), DeepSeek-R1-Distill-Qwen (1.5B/7B/32B), DeepSeek-R1-Distill-LLaMA-8B (Guo et al., 2025), QwQ-32B (Team, 2025), GLM-Z1-32B-0414 (GLM et al., 2024), s1-32B (Muennighoff et al., 2025), and DeepSeek-R1-671B (Guo et al., 2025). Among the closed-source models, we include Claude-3.5-Sonnet-1022, GPT-4o-0522, and OpenAI-o3-mini-medium.
At the small scale, MetaStone-S1-1.5B and MetaStone-S1-7B consistently outperform the listed open-sourced models with comparable or larger parameter sizes. Especially, MetaStone-S1-1.5B-low surpasses DeepScaleR-1.5B-Preview and R1-Distill-Qwen-1.5B on all datasets. And MetaStone-S1-1.5B-high further outperforms R1-Distill-Qwen-7B and R1-Distill-Llama-8B on AIME24 (57.9% vs 55.5%), demonstrating strong efficiency and capability of our lightweight SPRM. For MetaStone-S1-7B, its low reasoning effort has outperformed R1-Distill-Qwen-7B and R1-Distill-Llama-8B on AIME24 and LiveCodeBench. And MetaStone-S1-7B-high further gains the improvement of 14.7 points on AIME24 (70.2% vs 55.5%) and 4.8 points on LiveCodeBench (44.4% vs 39.6%).
At the larger scale, MetaStone-S1-32B-high achieves superior results on mathematical reasoning tasks, outperforming all listed open-sourced models of comparable or even larger size by +4.4% on AIME24 (85.2% vs 80.8%) and +3.6% on AIME25 (73.6% vs 70.0%). For other out-of-distribution tasks, MetaStone-S1-32B-high surpasses other 32B-sized models by +0.8% on LiveCodeBench (63.4% vs 64.2%) and +1.3% on C-EVAL (89.7% vs 88.4%). Compared to closed-source models, medium and high levels of MetaStone-S1-32B outperform Claude-3.5-Sonnet-1022 and GPT-4o-0522, and achieving comparable performance with medium level of OpenAI-o3-mini (85.2% vs 79.6% on AIME24, 73.6% vs 74.8% on AIME25, 64.2% vs 67.4% on LiveCodeBench, 89.7% vs 75.9% on C-EVAL,), highlighting its strong competitiveness in both mathematical and general reasoning tasks.
4.3 Ablation Study
Effectiveness of SPRM.
As shown in Table.2, our SPRM, with only around 26M extra parameters, achieves higher performance compared with 72B ORM and PRM. This demonstrates that sharing parameters between the reward model and policy model is able to generate high-quality guidance without the requirement of additional large-scale reward models. To further analyze the efficacy of different reward models, we compare their performance under different candidates $k$ in Fig.3. Across different $k$ and model sizes, SPRM consistently outperforms other methods, indicating its strong ability to distinguish between high and low quality reasoning trajectories.
Moreover, we evaluate the generalization ability of SPRM on tasks from the other domain (LiveCodeBench). Without any task-specific fine-tuning or in-domain data, our SPRM still achieves superior performance compared to separate reward models. This demonstrates SPRM’s strong zero-shot generalization capability, suggesting that it can capture domain-agnostic patterns to evaluate the reasoning trajectories.
To better show the effect of SPRM, in Fig.4, we visualize an example. The proposed SPRM correctly detect the calculation error at corresponding steps, results in a low score on both step and sample level.
Model Reward Model Extra Params AIME24 AIME25 LiveCodeBench
MetaStone-S1-1.5B-high Qwen2.5-Math-RM 72B 55.8 35.3 26.8
Qwen2.5-Math-PRM 72B 56.7 40.0 26.8
SPRM 5M 57.9 40.4 28.1
MetaStone-S1-7B-high Qwen2.5-Math-RM 72B 63.5 46.7 42.3
Qwen2.5-Math-PRM 72B 68.8 48.3 42.8
SPRM 26M 70.2 48.6 44.4
Table 2: Comparison of SPRM and other PRM models.
Figure 3: Evaluation of different number of candidates on AIME24.
[…]
Figure 4: SPRM predcitions on reasoning trajectory. Only key steps are listed. Correct and error steps are marked in green and red.
Effectiveness of self-supervised optimization.
We evaluate the effectiveness of SPRLoss in Table.3. Compared with using the final answer correctness as step-level supervision for PRM training, our proposed self-supervised optimization method achieves larger performance gains on both 1.5B and 7B models. Furthermore, Fig.5 shows the prediction score gap between correct and incorrect solutions. Compared to the BCELoss, SPRLoss demonstrates stronger discriminative capability with a larger score gap. This indicates that treating final answer correctness as step-level labels introduces substantial label noise, which harms the optimization. In contrast, SPRLoss leverages self-supervised signals to reduce the impact of noisy supervision, leading to more stable and accurate training.
Model Loss AIME24 LiveCodeBench
MetaStone-S1-1.5B-high BCELoss 56.7 27.9
SPRLoss 57.9 28.1
MetaStone-S1-7B-high BCELoss 69.1 43.9
SPRLoss 70.2 44.4
Table 3: Evaluation on SPRLoss.
4.4 Scaling Law of the Reflective Generative Model
In Fig. 6, we present the scaling law for reflective generative models, which shows the relationship between the total reasoning computation in TTS and the final performance. Following Snell et al. (2024), we define the computation budget $C$ as the product of the model’s parameter (B) and the total number of reasoning tokens (k): $C=Param_{policy}\times Token_{infer}$. Notably, when the total reasoning length is scaled to more than 32 times the baseline (e.g., Best-of-64), the performance improves slowly. Therefore, we mainly focus on TTS results up to Best-of-32 for each model scale. We observe that the final performance shows a positive correlation with the logarithm of the computation budget (the specific scaling factor depends on the baseline model architecture). This indicates that the final performance of our model can be enhanced by exponentially scaling on parameter size or the reasoning length.
Figure 5: The prediction score gap between correct and incorrect solutions. The blue curve shows the SPRLoss. The red curve shows the BCELoss.
Figure 6: The scaling law of Reflective Generative Models.
4.5 Aha Moment of SPRM
In Fig. 7, we present the predicted scores of SPRM for both correct and incorrect reasoning trajectories throughout the training process. During the initial phase of training, the optimization trajectories for all samples follow a similar trend, indicating that SPRM has not yet learned to distinguish between correct and incorrect reasoning trajectories. However, after a certain number of training steps (e.g., around 10/20/50 steps, 1280/2560/6400 samples for MetaStone-S1-1.5B/7B/32B, respectively), we observe a distinct "aha moment" point where the optimization trends of different reasoning trajectories begin to diverge. This suggests that the model is starting to judge the correctness based on the reasoning contents. With this aha moment, SPRM can be progressively refined through our proposed self-supervised SPRLoss, leading to a clear score gap between correct and incorrect reasoning trajectories, and enabling effective TTS.
Specially, we show an example in Fig.8. We fix the reasoning trajectory and use SPRM with different steps for scoring. In this case, the model mistakenly confuses $\varepsilon_{k}$ and $\varepsilon_{k+1}$, resulting in an incorrect solution. The SPRM fails to recognize the error before the aha moment, while the SPRM after the aha moment can correctly discriminate it.
Figure 7: The training process of SPRM. The blue and red curves denote the final prediction score on correct and incorrect reasoning trajectories. The green dashed line indicates the "aha moment".
Example: Aha Moment of SPRM Question: Let the sequence of rationals $x_{1},x_{2},\dots$ be defined such that $x_{1}=\frac{25}{11}$ and $[x_{k+1}=\frac{1}{3}\left(x_{k}+\frac{1}{x_{k}}-1\right)].$ $x_{2025}$ can be expressed as $\frac{m}{n}$ for relatively prime positive integers $m$ and $n$. Find the remainder when $m+n$ is divided by $1000$. Reasoning Trajectories: Alternatively, perhaps make a substitution. Let me define $\mu_{k}=1/\varepsilon_{k}$. Then, $1/\varepsilon_{k+1}=-1/\varepsilon_{k}+(8/9)/\varepsilon_{k}²$ Multiply both sides by $\varepsilon_{k}²$: $\varepsilon_{k}=-\varepsilon_{k}+(8/9)$ Thus, $\varepsilon_{k}+\varepsilon_{k+1}=8/9$ But no, not quitequestion and the correct answer should be 21+49=70) Before aha moment: Step score: 0.52; Final score: 0.54 After aha moment: Step score: 0.12; Final score: 0.36
Figure 8: Comparision of PRM predcitions before and after aha moment. Only key steps are listed. The error steps are marked in red.
4.6 Extend on MCTS
Since SPRM produces step-level evaluations, it can be naturally used in step-level search-based TTS methods such as Monte Carlo Tree Search (MCTS). In our setup, instead of performing full simulations to the end of a solution trajectory, we directly use SPRM to estimate the value of each node during the search. This enables a more efficient evaluation at each expansion step. In the expanding stage, we expand 4 children for the selected node and generate 1024 tokens in each child node. To balance the computation cost, we set the maximum number of tokens in the MCTS process from 0 (without MCTS) to 160k for each question.
Table.4 presents the performance of MCTS on the AIME24 dataset. By increasing the maximum number of searching tokens, the performance improved from 39.3 to 52.8, demonstrating the effectiveness of SPRM in providing high-quality step-level guidance in the reasoning stage. However, the performance of MCTS is still lower than the results under the Best-of-N strategy reported in Table 1. This is primarily due to the computational overhead inherent in tree-based search methods, which leads to incomplete searching in our setting. Nonetheless, the observed gains over the baseline validate the potential of our SPRM for integration with more advanced search-based inference methods.
Searching Tokens(k) 0 40 80 120 160
Accuracy(%) 39.3 48.8 50.0 51.7 52.8
Table 4: Performance of MetaStone-S1-1.5B with MCTS on AIME24.
5 Conclusion
In this work, we introduce MetaStone-S1, a novel Reflective Generative model that achieves competitive performance with OpenAI’s o3 model. In MetaStone-S1, we propose a self-supervised process reward model (SPRM) which provides fine-grained step-level guidance without requiring manual annotations. SPRM introduces only a lightweight prediction head and shares the majority of its parameters with the policy model, allowing on-policy joint optimization. The Reflective Generative models with SPRM naturally suitable for test time scaling, enabling performance improvement across both in-domain and out-of-domain benchmarks. We also summarized a scaling law between total thinking length and TTS performance. Furthermore, we find that our SPRM exhibits strong potential when integrated with search-based TTS methods, such as Monte Carlo Tree Search (MCTS). In future work, we will further explore its capability on step-level searching-based TTS and investigate its potential to enable real-time reasoning enhancement during inference.